OpenStudy (anonymous):

Solve for w. 5w + 9z = 2z + 3w

OpenStudy (anonymous):

This is in the wrong place but ill answer it subtract 3w from both sides and 2z-9z, you get 2w=-7z, so then divide both by -7 and you get (2w)/-7=z

OpenStudy (hockeychick23):

5w+9z-3w=2z+3w-3w 2w+9z-9z=2z-9z 2w= -7z 2w/2= -7/2z w=-7/2z